OpenStudy (anonymous):

Find the value(s) of a making v=7ai−2j parallel to w=a^2i+2j.

OpenStudy (anonymous):

Two vectors v and w are parallel if their cross product is zero

OpenStudy (anonymous):

So would you get the equation 7ai(a^2i)+(-2j)(2j) = 0 and then solve for a?

ganeshie8 (ganeshie8):

or, if the ratio of components are proportional

ganeshie8 (ganeshie8):

7a/a^2 = -2/2 solve a

OpenStudy (anonymous):

I solved that for -7 but I need a second answer

OpenStudy (anonymous):

i think 0 and -7

OpenStudy (anonymous):

Thank you I did not even think 0 would be an option.

ganeshie8 (ganeshie8):

me neither lol yes 0 is also an answer !

OpenStudy (anonymous):

a^2+7a=0 a(a+7)=0